OpenStudy (anonymous):

Y intercept = 50, 000 rate = +7% Y intercept = 70, 000 rate = -4% How can i find out When will these lines cross each other?

OpenStudy (kropot72):

If it is assumed that the rates correspond to slopes of 7/100 and -4/100 we can write two equations in the standard form y = mx + b as follows: \[y=\frac{7}{100}x+50000\ ...........(1)\] \[y=-\frac{4}{100}x+70000\ ...........(2)\] Subtracting equation (2) from equation (1) gives \[0=\frac{7}{100}x+\frac{4}{100}x-20000\ .............(3)\] Simplifying and rearranging equation (4) gives \[\frac{11}{100}x=20000\ ...........(4)\] Solving equation (4) will give the x coordinate of the solution and substitution of this value of x in either equation (1) or (2) will give the y coordinate of the solution